OpenStudy (anonymous):

Which is the equation of the line with slope 3 that contains point (-1, 5)? a)y – 1 = 3(x + 5) b)y – 5 = 3(x + 1) c)y – 1 = 3(x – 5) d)y – 5 = 3(x - 1)

OpenStudy (lncognlto):

Substitute the info you are given in the formula y - y1 = m(x - x1). What do you get?

OpenStudy (nurali):

y-y1=m(x-x1) (x1,y1)=(-1,5) m=3 y-5=3(x+1)